A string with a length of 3 m has two adjacent resonances at frequencies 112 Hz and 140 Hz. Determine the speed of the waves on the string vibrating at 140 Hz.

1
Expert's answer
2021-12-20T10:29:27-0500

Explanations & Calculations


  • You need to use the formula of the harmonics of this arrangement.
  • Let's assume the thread is fixed at both ends and only there.
  • The formula would be then,

"\qquad\qquad
\begin{aligned}
\small f_n&=\small\frac{nv}{2L}
\end{aligned}"

  • Let's take the consecutive harmonics at "n\,\&\,(n+1)" then,

"\qquad\qquad
\begin{aligned}
\small 112 &=\small \frac{nv}{2L}\implies n=\frac{112\times2L}{v}\cdots(1)\\
\\
\small 140&=\small\frac{(n+1)v}{2L}\implies n+1=\frac{140\times2L}{v}\cdots(2)\\
\\
\small n:(1)&\to(2),\\
\\
\small \frac{112\times2L}{v}+1&=\small \frac{140\times2L}{v}\\
\end{aligned}"

  • Now you can substitute for the length of the string and give it a try to get the answer.


  • A string in this arrangement can experience all the harmonics "\small n=1,2,3..."
  • But we do not know what those given frequencies exactly corresponds to and that we used algebraic notation for that.
  • The wave speed on a string remains constant as long as its tension, length and mass are kept unchanged.